what is the supplement of an angle that measures (40-2x)

Answer by jim_thompson5910(35256)

Supplementary Angle = 180 - (Given Angle)

Supplementary Angle = 180 - (40-2x)

Supplementary Angle = 180 - 40 + 2x

Supplementary Angle = 2x + 140